1. Gold (Au).
2. Titanium (Ti).
3. Gold/Beryllium (Au(99%) : Be(1%)).
Status of the evaporator - ( Operational)
The rotator doesn't work. Do not use rotation until further notice.
The updated operation procedure can be found here EDWARDS_SOP_updated_2.doc
, a laminated copy will be placed in the cleanroom. Talk to Oleh before using evaporator.
The operation procedure for the evaporator has changed. Do not use the evaporator
before talking to Oleh.
- The evaporator has been cleaned over the Xmas shutdown.
- Boats replaced.
- Note that for AuBe we use much smaller boat, that means you need to pass much less current (~3 A) to get to desired rate (~4 Å/s).
- AuBe pieces are large, you need to cut it with AuBe cutters and weigh using scales in the Fish tank lab. 330 mg gives ~180 nm. For 100 nm use ~230 mg.
- Run the evaporation till the boat is empty.
A new crystal monitor was obtained by Frank for this evaporator because the old crystal monitor kept on crashing during evaporation and Frank was unable to find out what was wrong with the old crystal monitor. However, the new crystal monitor made non-stop clicking noise when it was first plugged into the evaporator. Frank was unable to start up the crystal monitor properly and has decided to send the new crystal monitor back to US as it is still under warranty. Oleh has suggested to try an old crystal monitor that was used for the evaporator sitting in LG51. We are in the process of installing and trying out the crystal monitor from LG51.
The crystal monitor from LG51 was installed by Oleh and Jack. Oleh has done test evaporations and found that the rate showing on the crystal monitor was not very stable but it does work. We need to control the shutter and zero the thickness reading manually now.
Operation manual SQM160_Manual_V409.pdf